# Ricci-Notation Tensor Framework for Model-Based Approaches to Imaging

### 10 Jan 2024 v2

@@ IV Tensor Software
Table I introduces the RTToolbox. Together with the classes and functions of the MATLAB kernel, compiled or interpreted, the RTToolbox defines the RT software. First, this section offers a design overview of tensor objects. Next, it addresses index objects, which play a critical role in parsing a dual-variant index notation. Finally, key problems and solutions, contrasted with closest equivalents of the NT software, are summarized for unary, binary, and $N$-ary operations.
@@ IV-A Tensor Objects
Table II summarizes the tensor class. Apart from the first group, constructor aside, and isequal, all methods return a tensor object. Table III presents excerpts, cross-referenced to equations, from selected unit tests of class tensor.
@@
TABLE III: Tensor object unit tests. Included with the RTToolbox, the tensor object unit tests are chosen for their use-case value. To illustrate programmatic efficiency, software expressions are cross-referenced to equivalent algebraic expressions. Here, DN means any-degree and D# means degree-#.
@@
A tensor object has two protected properties, indices and entries, which the index and entry methods offer a copy of, respectively. The indices property is a row vector, possibly empty, of class index. The entries property must be a non-tensor array, \@iaciMDA MDA, usually of class double or logical, the only classes represented in unit tests.
@@ IV-B Index Objects
TABLE IV: Index object class definition. Having few of its own methods, the index class inherits pass-by-reference copying and (in)equality operators from the handle class. Inheritance enables use of set-theoretic functions exploited by the tensor class to parse an extended Ricci notation.
@@ IV-C Unary Operations
Unlike subsasgn, subsref never invokes permute. A subsref operation with index subscripts ignores preexisting indices of the tensor operand. Concatenated index subscripts simply redefine indices of the returned tensor, notwithstanding contractions and attractions, which eliminate indices, reduce degree, and modify entries.
@@ IV-D N-ary Operations
The alignn method inputs $N$ and outputs $N+2$ arguments. For each tensor input only its entries, \@iaciMDA MDA, is output. Via a kernel permute, the MDA is modified so that its dimensions align based on the set-theoretic union, in sequence, of indices irrespective of variants. A non-tensor input passes to an output only if it is a 2D array. One additional output is the set-theoretic union of indices, variants and left-to-right sequence preserved. If two variants of an index appear the leftmost is kept. The second additional output, an index vector, specifies which indices require summation, i.e., ones where two variants appeared.
@@
In theory, MATLAB could invoke a single overloaded method for an $N$-ary product expression involving a tensor. However, it invokes a sequence of binary product methods with a left-to-right precedence rule. Prior to RT software expression, this rule may require modification of \@iaciRT RT algebraic expression. In particular, to express an $N$-ary inner product over an index all but the rightmost operand must have the same variant, true or false, for that index, with only the rightmost operand having a complementary variant.
The NT software has two deviations from its NT algebra. To correctly compute some expressions, one or more sub-expressions have to be enclosed in parentheses and preceded by an extra operator, whose purpose is to decrement internal counters associated with underline-equivalent operators on enclosed indices. For this purpose, the tilde operator is hijacked. Thus, it cannot represent the entrywise NOT of a DenseNT, a tensor class of the NTToolbox. These deviations represent additional weaknesses, this time from a constructivism angle, of NT underline and multi-variant index formalisms.
@@ IV-E Binary Operations
Remaining binary methods, mtimes, mldivide, and mrdivide, are called pagewise because they exploit kernel pagemtimes, pagemldivide, and pagemrdivide functions through protected methods, namely mbinary, mbinary2, align2, and pageify. Each pagewise binary method invokes the mbinary method with a kernel function handle, e.g., @pagemtimes, as the first argument. For computational efficiency, the design involves a minimal number of kernel reshape and kernel permute invocations.
@@
As with scalar division, indices of denominator nonscalar operands have to change variants. Otherwise, the RT framework would be inconsistent in the context of linear systems. Prior to invoking mbinary, the mldivide and mrdivide methods complement indices using the not method.
The mbinary method receives the operands of a pagewise binary method as second and third arguments. If either operand is not a tensor then it must be a 2D array to avoid a runtime error by design. With a 2D array operand the tensor result has the same indices as the tensor operand. Resulting entries equal the result of a pagewise function, specified via a handle, applied to the tensor operand’s entries and the non-tensor operand. Because linear system operations are non-commutative, in general, pagewise function operands follow the original sequence received by mbinary.
When both operands of mbinary are tensor objects, the method invokes mbinary2 on them. After mbinary2 returns, mbinary has from it the final indices of the operation, two MDAs, and two vectors that specify a kernel reshape-and-permute. To compute the final entries of the original operation, mbinary applies a pagewise function, specified via a handle, to the two MDAs, in the order received, followed by the kernel reshape-and-permute.
The mbinary2 method invokes an align2 method that depends only on indices of each operand. This results in matching and leftover indices for each operand. Additional vectors indicate whether matching indices have the same variants. After align2 is done, mbinary2 knows which dimensions of each operand’s entries correspond in inner, outer, and entrywise fashion. Using this information, a kernel permute-and-reshape is applied separately, via the pageify method, to each operand’s entries. Along with resulting MDAs, the mbinary2 method returns information to mbinary for its kernel reshape-and-permute.
@@
Figure 7: Abstraction penalties vs. dimension size. The interpreted RTToolbox is computationally efficient when underlying unary, binary, or ternary operations involve large enough operands. In these tests, tensor operands have the same dimension sizes, $100\times 100\times P$. Tests vary a page dimension size, $P$.
@@
The RT software expresses the RT algebra with programmatic efficiency. When abstraction penalties are small, relative to equivalent computations requiring only the kernel and novel pagewise functions, the RT software achieves computational efficiency. Currently, kernel set-theoretic functions are interpreted not compiled. As they are critical to the RT abstraction, they are a key source of penalties. So that the RTToolbox obeys toolbox recommendations of MATLAB Central File Exchange, it was designed without compilable MEX files, such as C/C++ implementations of kernel set-theoretic functions.
@@ V Conclusions
This paper introduced \@iaciRT RT framework, the successor to a published NT framework, both of which complement a popular EMV framework. Comprising an underlying RT algebra and codesigned RT software, to model a problem and develop a solution, the proposed RT framework represents an approach for model-based imaging involving numeric tensors.
@@
The RT algebra inherits advantages of the NT algebra, with respect to $N$-degree support, associativity, commutativity, entrywise products, linear invertibility, and other published considerations. However, due to its dual-variant index notation, as opposed to the de facto multi-variant index notation of the NT algebra, the RT algebra provides all the expressivity of the NT algebra in a simpler way. Thanks to additional outer operations, inspired by EMV broadcasting, the RT algebra is also more expressive than the NT one. As it aligns well with the Ricci notation of geometric tensor calculus, the RT algebra lays a foundation for a deeper approach to imaging.
@@
Relevance of the RT algebra for model-based approaches to imaging is illustrated with an example, which demonstrates an asymptotically-efficient solution of an optimization problem where the SSE, a scalar function of deviant pixels in the image plane, depends on an unknown phase aberration in the pupil plane, a Fourier domain. The RT algebra helps to model and simplify the SSE and its first and second derivatives, namely a gradient and \@iaciHMF HMF, where the complexity of computing all three proves equivalent to computing just the SSE only. The approach makes visible diffraction rings and exoplanets near an occulted star in a simulated coronagraph image.
@@
Thanks to the RT algebra, all nonzero-degree matrices in the gradient expression were eliminated. Moreover, gradient and HMF expressions were reorganized to reveal embedded 2D (inverse) DFTs, acceleratable via 2D (inverse) FFTs. Whereas the asymptotic time-and-space efficiencies were presented using MATLAB without the RTToolbox, the latter helped. A natural implementation of the RT algebra simplified a validation of intermediate steps in the underlying derivation. During the search for asymptotic efficiencies, algebraic errors were corrected via related RT software sanity checks.
@@
The RT software is defined as the RTToolbox, comprising tensor and index classes, new pagewise functions, and unit tests, plus MATLAB. In this manner, the RT framework complements a popular EMV framework. Methods, functions, and tests presented, a subset of what could be developed, are chosen to emphasize key software requirements. Arguing that additional computational efficiency is possible, the approach taken emphasizes programmatic efficiency. Allowing the RT algebra to be naturally expressed with MATLAB, the RTToolbox is designed in a way where small abstraction penalties, related to index objects, would reduce with a MATLAB release that compiled kernel set-theoretic functions.
@@
Compared to the NT software for MATLAB end users, comprising a C/C++ library called LibNT, M-file and MEX wrappers of the NTToolbox, plus MATLAB itself, the RT software is simpler, comprising only M-files of the RTToolbox plus MATLAB. The toolbox uses an object-oriented approach to realize a dual-variant index notation, with its index class inheriting a kernel handle class, congruous to C/C++ void pointers, for extra programmatic and computational efficiency. While the NTToolbox has a tensor class called DenseNT, it relies on MATLAB character vectors and string processing to represent and parse a multi-variant index notation.
@@
Finally, the paper elaborated on unary, binary, and $N$-ary tensor operations. For example, to express an $N$-ary inner product over an index, the RT software may vary from the RT algebra. As MATLAB parses an $N$-ary product into a binary product sequence with left-to-right precedence, all but the rightmost one must be entrywise to express the $N$-ary inner product. This variation is contrasted with those of the NT software, which unlike the NT algebra required some products to be enclosed in parentheses and preceded with an additional operator. Other than introducing a left-to-right precedence rule, the RT software expresses the RT algebra faithfully.
@@ Acknowledgments
The author gratefully acknowledges editorial advice from Dan Sirbu, who suggested motivating the RT framework with a model-based example involving DFTs, and Adam P. Harrison, whose NT framework laid a foundation for this work.
Code and Data Availability
Code is shared with a permissive free software licence. For the imaging example, including its source image, visit www.github.com/KoderKong/Fourier2D. For the RTToolbox, visit www.mathworks.com/matlabcentral/fileexchange/156174-ricci-notation-tensor-toolbox-rttoolbox.

### 08 Apr 2024 v3

@@ IV Tensor Software
Table I introduces the RTToolbox. Together with the classes and functions of the MATLAB kernel, compiled or interpreted, the RTToolbox defines the RT software. First, this section offers a design overview of tensor objects. Next, it presents index objects, which play a critical role in parsing a dual-variant index notation. Finally, key problems and solutions, contrasted with closest equivalents of the NT software, are summarized for unary, binary, and $N$-ary operations.
@@ IV-A Tensor Objects
Table II summarizes the tensor class. Apart from the first group, constructor aside, and isequal, all methods return a tensor object. Table III presents excerpts, cross-referenced to selected equations, from unit tests of class tensor.
@@
TABLE III: Tensor object unit tests. Included with the RTToolbox, the tensor object unit tests were developed for this paper. To illustrate programmatic efficiency, software expressions are cross-referenced to equivalent algebraic expressions. Here, DN means any degree and D# means degree #.
@@
A tensor object has two protected properties, indices and entries, which the index and entry methods offer a copy of, respectively. The indices property is a row vector, possibly empty, of class index. The entries property must be a non-tensor array, an MDA, usually of class double or logical, the only classes represented in unit tests.
@@ IV-B Index Objects
TABLE IV: Index object class definition. Having few of its own methods, the index class inherits copy-by-reference, (in)equality, and concatenation operators from the handle class. Inheritance enables use of set-theoretic functions exploited by the tensor class to parse an extended Ricci notation.
@@ IV-C Unary Operations
Unlike subsasgn, subsref never invokes permute. A subsref operation with index subscripts ignores preexisting indices of the tensor operand. Concatenated index subscripts simply redefine indices of the returned tensor, notwithstanding contractions and attractions, which simplify entries and shorten indices, reducing degree.
@@ IV-D N-ary Operations
The alignn method inputs $N$ and outputs $N+2$ arguments. For each tensor input only its entries, an MDA, is output. Via a kernel permute, the MDA is modified so that its dimensions align based on the set-theoretic union, in sequence, of indices irrespective of variants. A non-tensor input passes to an output only if it is a 2D array. One additional output is the set-theoretic union of indices, variants and left-to-right sequence preserved. If two variants of an index appear the leftmost is kept. The second additional output, an index vector, specifies which indices require summation, i.e., ones where two variants appeared.
@@
In theory, MATLAB could invoke a single overloaded method for an $N$-ary product expression involving a tensor. However, it invokes a sequence of binary product methods with a left-to-right precedence rule. Prior to RT software expression, this rule may require modification of an RT algebraic expression. In particular, to express an $N$-ary inner product over an index all but the rightmost operand must have the same variant, true or false, for that index, with only the rightmost operand having a complementary variant.
The NT software has two deviations from its NT algebra. To correctly compute some expressions, one or more sub-expressions have to be enclosed in parentheses and preceded by an extra operator, whose purpose is to decrement internal counters associated with underline-equivalent operators on enclosed indices. For this purpose, the tilde operator is hijacked. Thus, it cannot represent the entrywise NOT of a DenseNT, a tensor class of the NT software. These deviations represent additional weaknesses, this time from a constructivism angle, of NT underline and multi-variant index formalisms.
@@ IV-E Binary Operations
Remaining binary methods, mtimes, mldivide, and mrdivide, are called pagewise because they exploit kernel pagemtimes, pagemldivide, and pagemrdivide functions through protected methods, namely mbinary, mbinary2, align2, and lattice. Each pagewise binary method invokes the mbinary method with a kernel function handle, e.g., @pagemtimes, as the first argument. For computational efficiency, the design involves a minimal number of kernel reshape and kernel permute invocations.
@@
As with scalar division, indices of denominator nonscalar operands have to change variants. Otherwise, the RT framework would be inconsistent in the context of linear systems. Prior to invoking mbinary, the mldivide and mrdivide methods complement indices via a not operation.
The mbinary method receives the operands of a pagewise binary method as second and third arguments. If either operand is not a tensor then it must be a 2D array to avoid a runtime error by design. With a 2D array operand the tensor result has the same indices as the tensor operand. Resulting entries equal the result of a pagewise function, specified by handle, applied to the tensor operand’s entries and the non-tensor operand. Because linear system operations are noncommutative, in general, pagewise function operands follow the original sequence received by mbinary.
When both operands of mbinary are tensor objects, the method invokes mbinary2 on them. After mbinary2 returns, mbinary gets from it the final indices of the operation, two MDAs, and two vectors that specify a kernel reshape-and-permute. To compute the final entries of the original operation, mbinary applies a pagewise function, specified via a handle, to the two MDAs, in the order received, followed by the kernel reshape-and-permute.
The mbinary2 method invokes an align2 method that depends only on indices of each operand. This results in matching and leftover indices for each operand. Additional vectors indicate whether matching indices have the same variants. After align2 is done, mbinary2 knows which dimensions of each operand’s entries correspond in inner, outer, and entrywise fashion. Using this information, a kernel permute-and-reshape is applied separately, via the lattice method, to each operand’s entries. Along with resulting MDAs, the mbinary2 method returns information to mbinary for its kernel reshape-and-permute.
@@
Figure 7: Abstraction penalties vs. dimension size. The interpreted RTToolbox is computationally efficient when underlying unary, binary, or ternary operations involve large enough operands. In these tests, tensor operands have the same dimension sizes, $100\times 100\times P$. Tests vary just one dimension size, $P$.
@@
The RT software expresses the RT algebra with programmatic efficiency. When abstraction penalties are small, relative to equivalent computations requiring only the kernel and novel pagewise functions, the RT software achieves computational efficiency. Because kernel ismember and unique functions, critical to parsing the RT algebra, are interpreted not compiled, they are a key source of penalties. So that the RTToolbox obeys toolbox recommendations of MATLAB Central File Exchange, it was designed without compilable MEX files, such as C/C++ implementations of kernel set-theoretic functions.
@@ V Conclusions
This paper introduced an RT framework, the successor to a published NT framework, both of which complement a popular EMV framework. Comprising an underlying RT algebra and codesigned RT software, to model a problem and develop a solution, the proposed RT framework represents an approach for model-based imaging involving numeric tensors.
@@
The RT algebra inherits advantages of the NT algebra, with respect to $N$-degree support, associativity, commutativity, entrywise products, linear invertibility, and other published considerations. However, due to its dual-variant index notation, as opposed to the de facto multi-variant index notation of the NT algebra, the RT algebra provides all the expressivity of the NT algebra in a simpler way. Thanks to additional outer operations, inspired by EMV broadcasting, the RT algebra is also more expressive than the NT one. Although it aligns well with the Ricci notation of geometric tensor calculus, this paper proposes the RT algebra for numeric tensor purposes.
@@
Relevance of the RT algebra for model-based approaches to imaging is illustrated with an example, which demonstrates an asymptotically-efficient solution of an optimization problem where the SSE, a scalar function of deviant pixels in the image plane, depends on an unknown phase aberration in the pupil plane, a Fourier domain. The RT algebra helps to model and simplify the SSE and its first and second derivatives, namely a gradient and an HMF, where the complexity of computing all three proves equivalent to computing just the SSE alone. The approach makes visible diffraction rings and exoplanets near an occulted star in a simulated coronagraph image.
@@
Thanks to the RT algebra, all nonzero-degree matrices in a gradient formulation were eliminated. Moreover, gradient and HMF expressions were reorganized to reveal (inverse) 2D DFTs, acceleratable via (inverse) 2D FFTs. Whereas the asymptotic time-and-space efficiencies were presented using MATLAB and not the RT software, the latter proved useful. A natural implementation of the RT algebra simplified a validation of intermediate steps in the underlying derivation. During the search for asymptotic efficiencies, algebraic errors were corrected via related RT software sanity checks.
@@
The RT software is defined as the RTToolbox, comprising tensor and index classes, new pagewise functions, and unit tests, plus MATLAB. In this manner, the RT framework complements a popular EMV framework. Methods, functions, and tests presented, a subset of what could be developed, were chosen to emphasize key software requirements. Arguing that additional computational efficiency is possible, the approach taken emphasizes programmatic efficiency. Allowing the RT algebra to be naturally expressed with MATLAB, the RTToolbox is designed in a way where small abstraction penalties, related to index objects, would reduce with a MATLAB release that compiled kernel set-theoretic functions.
@@
Compared to the NT software for MATLAB end users, comprising a C/C++ library called LibNT, M-file and MEX wrappers of the NTToolbox, plus MATLAB itself, the RT software is simpler, comprising only M-files of the RTToolbox plus MATLAB. The toolbox uses an object-oriented approach to realize a dual-variant index notation, with its index class inheriting a kernel handle class, congruous to C/C++ void pointers, for extra programmatic and computational efficiency. In contrast, while the NTToolbox has a tensor class called DenseNT, it relies on MATLAB character vectors and string processing to realize a multi-variant index notation.
@@
Finally, the paper elaborated on unary, binary, and $N$-ary tensor operations. For example, to express an $N$-ary inner product over an index, the RT software may vary from the RT algebra. As MATLAB parses an $N$-ary product into a binary product sequence with left-to-right precedence, all but the rightmost index must express an entrywise, not inner, product. This variation is contrasted with those of the NT software, which unlike the NT algebra required some products to be enclosed in parentheses and preceded with an additional operator. Other than introducing a left-to-right precedence rule, the RT software expresses the RT algebra faithfully.
@@ Acknowledgments
The author gratefully acknowledges editorial advice from Dan Sirbu, who suggested illustrating the RT framework with a model-based example involving DFTs, and Adam P. Harrison, whose NT framework laid a foundation for this work. He also thanks anonymous reviewers for constructive feedback that led to significant improvements, especially to clarity of the paper’s topics and goals and to its imaging example.
Code and Data
Code is shared with a permissive free software licence. For the imaging example, including its free source image, visit www.github.com/KoderKong/Fourier2D. For the RTToolbox, visit www.mathworks.com/matlabcentral/fileexchange/156174-ricci-notation-tensor-toolbox-rttoolbox.